# AdaptBot: Combining LLM with Knowledge Graphs and Human Input for Generic-to-Specific Task Decomposition and Knowledge Refinement

### v1

@@ Abstract
Embodied agents assisting humans are often asked to complete a new task in a new scenario. An agent preparing a particular dish in the kitchen based on a known recipe may be asked to prepare a new dish or to perform cleaning tasks in the storeroom. There may not be sufficient resources, e.g., time or labeled examples, to train the agent for these new situations. Large Language Models (LLMs) trained on considerable knowledge across many domains are able to predict a sequence of abstract actions for such new tasks and scenarios, although it may not be possible for the agent to execute this action sequence due to task-, agent-, or domain-specific constraints. Our framework addresses these challenges by leveraging the generic predictions provided by LLM and the prior domain-specific knowledge encoded in a Knowledge Graph (KG), enabling an agent to quickly adapt to new tasks and scenarios. The robot also solicits and uses human input as needed to refine its existing knowledge. Based on experimental evaluation over cooking and cleaning tasks in simulation domains, we demonstrate that the interplay between LLM, KG, and human input leads to substantial performance gains compared with just using the LLM output.
@@ 3.4 Knowledge refinement with human input
Existence check: if an action or object in the LLM output does not exist in the KG, there are three possibilities: (1) The agent is mistaking an existing item (action) for another item (action); (2) the entity does not exist in the domain; or (3) the entity exists but is not in the KG. In the first case, human informs the agent about the correct object (or action); in the second case, human denies existence of entity; and in the third case, human confirms the entity’s existence and agent interactively obtains entity’s attributesSupplementary material includes details of questions asked..
@@ 4.1.4 Evaluation Strategy
For the evaluation of our framework, we used human participants to provide ground truth. Specifically, we recruited 18 human evaluators to mark the execution outputs for each task assigned to the framework and the two baselines. The tasks were distributed such that the output for each task was evaluated by at least three human participants. The scores provided by the human (on a linear scale between 0-20) were averaged to obtain the success rate of our framework and the two baselines. These results are discussed further in Section 4.2.
@@
4.1.5 Evaluation Measures
The key performance measures considered in this work include:
4.2 Experimental Results
Next, we describe and discuss the experimental results.
@@ 4.2 Experimental Results
Our project web sitehttps://sssshivvvv.github.io/adaptbot/ hosts our supplementary material, including examples of tasks being performed in simulation, and supporting results with other LLM models.

### v2

@@ Abstract
An embodied agent assisting humans is often asked to complete new tasks, and there may not be sufficient time or labeled examples to train the agent to perform these new tasks. Large Language Models (LLMs) trained on considerable knowledge across many domains can be used to predict a sequence of abstract actions for completing such tasks, although the agent may not be able to execute this sequence due to task-, agent-, or domain-specific constraints. Our framework addresses these challenges by leveraging the generic predictions provided by LLM and the prior domain knowledge encoded in a Knowledge Graph (KG), enabling an agent to quickly adapt to new tasks. The robot also solicits and uses human input as needed to refine its existing knowledge. Based on experimental evaluation in the context of cooking and cleaning tasks in simulation domains, we demonstrate that the interplay between LLM, KG, and human input leads to substantial performance gains compared with just using the LLM.
@@ 3.4 Knowledge refinement with human input
Existence check: if an action or object in the LLM output does not exist in the KG, there are three possibilities: (1) The agent is mistaking an existing item (action) for another item (action); (2) the entity does not exist in the domain; or (3) the entity exists but is not in the KG. In the first case, human informs the agent about the correct object (or action); in the second case, human denies existence of entity; and in the third case, human confirms the entity’s existence and agent interactively obtains entity’s attributesSupplementary material includes details of questions asked.. As the agent expands its knowledge, the need for human input progressively decreases.
@@ 4.1.4 Evaluation Strategy
For the evaluation of our framework, we used human participants to provide ground truth. Specifically, we recruited 18 human evaluators to mark the execution outputs for each task assigned to the framework and the two baselines. These evaluators were subject matter experts who were not involved in the design of our framework. The tasks were distributed such that the output for each task was evaluated by at least three human participants. The scores provided by the human (on a linear scale between 0-20) were averaged to obtain the success rate of our framework and the two baselines. These results are discussed further in Section 4.2.
@@
4.1.5 Evaluation Measures
The key performance measures considered in this work include:
4.2 Experimental Results
Next, we describe and discuss the experimental results.